# Manifesto from Dagstuhl Perspectives Workshop 24352—Conversational Agents: A Framework for Evaluation (CAFE)

### v1

@@ 4 Stakeholder Goals
Stakeholders have varying goals, which are not generally accessible to system designers or evaluators and are not always explicitly formulated even for the stakeholder. They can often be inferred from the actions and the background information the system has available. Stakeholder goals are not always aligned: a system may benefit some stakeholders more than others [112, 39]. In general, stakeholders may have several simultaneous, interacting, and only partially overlapping goals [38, 12]. A conversational approach can further many stakeholder goals.
  • Users have goals they wish to see a system help them achieve. A user may be solving an immediate problem or a long-term issue; wish to retrieve something they already know about (i.e., known item search); seek to learn or investigate a topical or thematic area [84]; they may be looking for diversion or entertainment or delight. Users may be focused on a goal, or the goal may be in the background for some other activity, and the goal may be of varying importance and centrality to the user throughout the task they are pursuing. Users may approach information access systems to find immediate factual responses, seek support for an argument, or plan a later investigation into an area. They may be seeking documents or collections of documents, they may be looking for some factual content they know or believe to be found in a document, or they may be looking to enjoy reading, viewing, or listening to some content that may be found in an item in a collection or across several items. To achieve these goals, users will generally want to expend little effort, enjoy interacting with the system, trust the correctness of the result, and feel confident that they have exhausted the usefulness of the system they are working with for the purposes they have engaged with it.
@@
As deeply discussed in many venues [2], relevance, truthfulness, and trustworthiness [^3] https://digital-strategy.ec.europa.eu/en/library/ethics-guidelines-trustworthy-ai are primary and very common top-level goals. These top-level goals are typically well-represented and adequately considered in many evaluation activities in the field. If you look at the tasks offered by initiatives such as Text REtrieval Conference (TREC) [50], Conference and Labs of the Evaluation Forum (CLEF) [42, 41], NII Testbeds and Community for Information access Research (NTCIR) [98], or Forum for Information Retrieval Evaluation (FIRE), you will find many evaluation efforts revolving around these top-level goals.
@@
Users wish to minimize the effort they expend on using a system. This means the interaction must have low friction and a low threshold. A conversational system will allow users to approach the system with little preparation and low cognitive load: they do not need to formulate their needs, tasks, and goals in terms exactly acceptable to the system since this can be negotiated during the session [117]. Conversely, using a conversational system may incur overhead—instead of immediate access to a known item through a short command, a user may need to engage in a more verbose interaction. A system needs to be convenient to the frequent power user and forgiving to the novice and the occasional user. The latter has more to gain from a conversational approach.
@@
This allows users to specialize their goals over the course of a conversation, starting with a general exploration of a topical area or a theme and then progressively drilling down into a more detailed or fine-grained view until some target of interest is found. Iterative and interactive information retrieval provides a basis to understand some of the challenges, most notably that of query reformulation, which, without the support of a system, has been known to be a demanding task for users [14].
Alternatively, this allows users find and pursue a trajectory or a curriculum through a collection to meet a less specific and possibly longer-term goal of learning or familiarizing oneself with a topic. Achieving expertise in a topic does not only involve having access to documents and materials but also acquiring skills that are imparted through conversation and apprenticeship with an expert [29, 96].
@@
Among several conversational goals are those of establishing a relationship between counterparts. This is an incidental goal that in human-human interaction is addressed in parallel with informational or narrational activities [57, 24], which has been studied in human-human communication [124].
@@ 5 User Aspects
Some users have preferences, special needs, or other characteristics that will make a conversational system more useful than a non-conversational one; for example, people with visual impairments, senior citizens, children, or those seeking coaching. Following principles of Universal Design [47, 87] or “Design for All”, many or even most such specific preferences will lead to design solutions that improve interaction for every user in the general case. Contextual requirements may benefit from a conversational approach: hands-free operation of a system or situations where a user’s visual attention is otherwise occupied are examples where a conversational system would well serve every user.
@@
Table 1: A model of health information seekers, based on Pang et al. [89]
@@
Users also differ in their cognitive needs around information access, e.g., in their tendency to seek information. For example, the health domain Pang et al. [89] categorizes information seekers based on their level of reading engagement and research tactics: a knowledge digger would be high on both, whereas a quick fact checker would be low on both. In decision-making, a similar distinction is made between users who want to make the best, most optimal decision possible (maximizers) versus those who just want a satisfactory outcome (satisfiers) that satisfies most of their needs. Maximizers usually engage in extensive product comparison and aspire to find the best possible option [27]. In earlier work on a critiquing recommender system, Frederix [43] found that maximizers experienced more guidance from a (conversational) critiquing system compared to a simple filtering tool, whereas novices did not. Both these examples show that conversational systems might be used, experienced, and thus evaluated quite differently depending on individual differences of the user. Using individual difference measures (such as a maximization scale) allows one to distinguish between users in the evaluation.
@@
Cultural differences in how a conversation proceeds, if not accommodated by a conversational system, may annoy and disturb users [69]. For instance, users from cultures that value directness and efficiency might prefer concise, straightforward interactions, whereas those from cultures that emphasize politeness and formality might expect more elaborate and courteous exchanges. This can influence how questions are asked, responses are given, and even the pacing of the conversation. The notion of politeness in itself is a many-faceted behavioral characteristic, and its various aspects are prioritized differently across cultures.
@@ 6 Task
  • an information need and knowledge gap [35]
@@
  • interactive and mixed initiative [54]
@@
In contrast, some tasks strongly benefit from the opportunities afforded by conversation, including an iterative process and agents interacting via turn-taking. Particularly, complex tasks [19, 21] that require clarification and refinement to support a user’s goal are well-served by an interaction between the system and the user.
@@
Figure 7: The tasks that are ideally solved by CONIAC systems. The model proposes two core dimensions of task complexity and task definiteness. (Iterative vs. exploratory search behavior representation adapted from White and Roth [123])
In this model, complex tasks can vary in definiteness, with implications for the nature of an interaction between a system and a user. A complex task that is more well-defined requires primarily an iterative interaction process aimed at refining the expression of the information need to be more focused or precise [14]. A more ill-defined information need involves exploration, a process that involves learning and investigation, and examining a broader range of information—often uncovering unanticipated information. This process typically involves refining the information need itself. Here, we draw on the notion of iterative vs. exploratory search defined by White and Roth [123]. This model explicitly reflects the interaction behaviors that are likely to be involved with these two types of search.
A user’s information journey in complex tasks is well-documented in the literature [84, 74], highlighting how user states can evolve throughout the process. For instance, the Information Search Process (ISP) model [74] demonstrates that a user’s initial perception of a task’s complexity might start as simple but can quickly become more complex due to the broad scope of relevant information. This suggests that CONIAC systems should be capable of detecting and modeling these evolving user states, optimizing their functions and responses accordingly to enhance the user experience.
Task switching is another common characteristic of human information-seeking behavior, observed in both web search engines and conversational assistants [108, 119]. In complex tasks, users often pursue multiple sub-goals, and the outcome of each sub-goal can influence subsequent actions or significantly impact the final goal. Therefore, it is crucial for users to track multiple tasks effectively. However, traditional information access systems typically leave the burden of managing these tasks to the users.
@@
In the following, we illustrate typical conversation information access tasks and how they relate to dimensions of complexity and definiteness. We also illustrate how users might dynamically traverse the space during the conversational interaction with the system [114, 74].
@@
In the context of scientific discovery, hypothesis generation involves investigating scientific literature. This often includes making connections between seemingly unrelated studies [113]. A classic example is the “ABC” literature-based discovery model introduced by Swanson [106], in which implicit shared items (“B” items) are analyzed to form links between two entities (“A”, “C”). Under this model, a user could be supported by a conversational system to move logically between different parts of the information space.
Evaluation of a hypothesis generation task is challenging as there is no truly “correct” answer that the user seeks. Notions of novelty, surprise, feasibility, promisingness, or interestingness are relevant [105].
@@
Suppose a user with limited domain knowledge needs to replace an old washing machine: “I need to replace my washing machine and I want one that serves our family of five and is somewhat efficient”. We initially would classify this as a well-defined need of moderate complexity. Current online tools (e-commerce shops) would allow you to search products with some filtering and construct a set of alternatives, but it will be hard to understand the domain: e.g., what is a sufficient capacity in kg? What amount of kWh makes it efficient? Also, the number of features presented by the tool might bring the realization that there are other important things to consider (rinsing quality, time per cycle) and might overwhelm the user with the number of tradeoffs to make. Hence, a simple search might make the user aware that the need is not as well-defined and the task perhaps more complex than envisioned. However, a regular online shop will not provide the tools to help users understand the domain better. A conversational system, using techniques from critiquing recommenders [43] would help the user to understand the domain and tradeoffs better: “based on your query, we have several efficient washing machines that have 8kg capacity or more that suit your family needs. Would you rather like a washing machine that rinses the clothes better and but will be less silent”. Based on this response, the user discovers some other (latent) needs and tradeoffs they were unaware of before, and the dialog will allow them to define them better iteratively. This will allow the user to learn about their actual preferences and the best matching products, and the system can provide a more adequate recommendation.
@@
Looking for health information is a natural task for CONIAC. Many individuals already use the internet for health advice, such as using symptom checkers to determine if they might have an underlying condition or whether they should consult a healthcare professional [30].
Research has examined health information-seeking behaviors, for example, considering how users may address an open-ended task such as “Imagine you are going to a party and will discuss health information with your friends. Gather enough information for an interesting discussion.” (adapted from [90]). This is a vague but complex task, requiring exploration of a large and dynamic information space, accommodation of serendipity to access novel health concepts, and interactions between the user and the system to support the identification of information that is interesting and engaging to the user.
A recent study on a CONIAC system designed to help patients find cancer-related clinical trials indicates that these systems could make health information more accessible for individuals with limited health or computer literacy skills [17].
While CONIAC has significant potential, several concerns must be addressed when implementing these systems in the health domain. One issue is that these systems may lack the necessary expertise to accurately answer all questions, which could lead to misunderstandings or misinterpretations and, consequently, incorrect responses [111]. Although this is a common challenge for all search systems, it could be particularly catastrophic with health information. Additionally, these systems often handle sensitive patient data, which requires robust safeguarding measures. Voice-only CONIAC systems might also face difficulties with speech recognition, especially when users are distressed or in noisy environments [107].
@@
Enterprise information access and personal information access have distinct requirements compared to traditional web search engines. These include challenges such as searching across enterprise (individual) intranets or navigating multiple internal or private data sources [51]. Although there has been growing interest in workplace-oriented digital assistants, such as Alexa for Business or the Cortana Skills Kit for Enterprise, their adoption has been limited. However, with the increased adoption of LLM-based systems, new solutions like Microsoft Copilot and Google Gemini are emerging to enhance productivity and streamline information retrieval in enterprise settings [118]. Microsoft Copilot is designed to assist employees by integrating with Office applications and providing contextual assistance and insights with LLMs. Similarly, Google Gemini aims to leverage artificial intelligence to improve search and information access across Google’s ecosystem. Despite these advancements, the use of these tools remains relatively low, suggesting that further development is needed to realize and understand the possible task and the system’s potential in the enterprise or personal information environment fully [118, 119].
@@
Children regularly engage in information discovery tasks that involve exploring online resources to complete classroom assignments. Their default approach is to use web search engines, but they often face challenges: (i) difficulty in formulating their information needs into succinct keyword queries, (ii) a preference for using natural language or question-based queries, and (iii) trouble navigating Search Engine Result Pages to find relevant resources [9, 76]. A conversational information access system could be highly beneficial in this context. Such a system could ask clarifying questions to help children refine their information needs, provide specific answers when required, and recommend a limited set of resources for further exploration. However, a one-dimensional evaluation (e.g., the system’s ’goodness’ based on its ability to deliver the necessary information or user perception—especially, since children’s perceptions of their interactions with information systems do not always align with their actions) is insufficient. As children are the primary stakeholders in this scenario, it is essential to consider factors like the suitability of the information provided and whether the vocabulary matches their literacy levels. This highlights the importance of explicitly considering the task at hand, the stakeholder performing the task, and the various criteria for evaluation before determining the specific metrics to be used as we reflect on the CAFE components.
@@
“The next two and a half hours, I will be working on something that takes up most of my attention, gaze, and hands. I want to have something entertaining in my headphones during that time. I am interested in some general topics: personal aviation, allergy medication, history of Central Asia, and I like yacht rock and string quartets” (the last part is presumably known to the system). The system can suggest some listening material to me, excerpts from audiobooks and lectures, recent newscasts with local news, and playlists I have saved in the library; I can respond by ranking the suggestions, helping the system to sort the items that will be played in the session. The system can verify if those preferences are generally true or specific to this session.
@@
I ask a system about interesting sights to see in Saarland and bike paths. I get some tips, but they do not seem to be geared toward my tastes, so I elaborate that I want bike paths for mountain biking and that I will be camping. The system responds but delivers the same content. I ask for tips on Limburg, an area I have already vacationed in and know well; I realize that the knowledge of this system is superficial while it was initially presented as a useful resource, and I will now turn to another system. (If not conversational, I would have had to work hard to redo the query for another area)
@@
Figure 8: Visualization of possible tasks for information access, demonstrating different information access and activity levels [122, 120]. The diagram highlights the progression from simple information finding to more complex tasks like learning, investigating, critiquing, evaluating, extracting, and creating. The new frontier with Generative Information Retrieval (marked by dashes) indicates that these systems can enable advanced tasks such as critiquing and evaluating, expanding beyond the traditional search frontier also applicable to CONIAC tasks.

### v2

@@ 4 Stakeholder Goals
Stakeholders have varying goals, which are not generally accessible to system designers or evaluators and are not always explicitly formulated even for the stakeholder. They can often be inferred from the actions and the background information the system has available. Stakeholder goals are not always aligned: a system may benefit some stakeholders more than others [surer2018multistakeholder, ekstrand2024not]. In general, stakeholders may have several simultaneous, interacting, and only partially overlapping goals [INR-079, bauer2019_multimethod_multistakeholder]. A conversational approach can further many stakeholder goals.
  • Users have goals they wish to see a system help them achieve. A user may be solving an immediate problem or a long-term issue; wish to retrieve something they already know about (i.e., known item search); seek to learn or investigate a topical or thematic area [marchionini2006]; they may be looking for diversion or entertainment or delight. Users may be focused on a goal, or the goal may be in the background for some other activity, and the goal may be of varying importance and centrality to the user throughout the task they are pursuing. Users may approach information access systems to find immediate factual responses, seek support for an argument, or plan a later investigation into an area. They may be seeking documents or collections of documents, they may be looking for some factual content they know or believe to be found in a document, or they may be looking to enjoy reading, viewing, or listening to some content that may be found in an item in a collection or across several items. To achieve these goals, users will generally want to expend little effort, enjoy interacting with the system, trust the correctness of the result, and feel confident that they have exhausted the usefulness of the system they are working with for the purposes they have engaged with it.
@@
As deeply discussed in many venues [hleg2019ethics], relevance, truthfulness, and trustworthiness [^3] https://digital-strategy.ec.europa.eu/en/library/ethics-guidelines-trustworthy-ai are primary and very common top-level goals. These top-level goals are typically well-represented and adequately considered in many evaluation activities in the field. If you look at the tasks offered by initiatives such as Text REtrieval Conference (TREC) [zz-HarmanVoorhees2005-editor], Conference and Labs of the Evaluation Forum (CLEF) [zz-FerroPeters2019-editor, Ferro2024], NII Testbeds and Community for Information access Research (NTCIR) [zz-SakaiEtAl2020-editor], or Forum for Information Retrieval Evaluation (FIRE), you will find many evaluation efforts revolving around these top-level goals.
@@
Users wish to minimize the effort they expend on using a system. This means the interaction must have low friction and a low threshold. A conversational system will allow users to approach the system with little preparation and low cognitive load: they do not need to formulate their needs, tasks, and goals in terms exactly acceptable to the system since this can be negotiated during the session [trippas2020towards]. Conversely, using a conversational system may incur overhead—instead of immediate access to a known item through a short command, a user may need to engage in a more verbose interaction. A system needs to be convenient to the frequent power user and forgiving to the novice and the occasional user. The latter has more to gain from a conversational approach.
@@
This allows users to specialize their goals over the course of a conversation, starting with a general exploration of a topical area or a theme and then progressively drilling down into a more detailed or fine-grained view until some target of interest is found. Iterative and interactive information retrieval provides a basis to understand some of the challenges, most notably that of query reformulation, which, without the support of a system, has been known to be a demanding task for users [belkin2001iterative].
Alternatively, this allows users find and pursue a trajectory or a curriculum through a collection to meet a less specific and possibly longer-term goal of learning or familiarizing oneself with a topic. Achieving expertise in a topic does not only involve having access to documents and materials but also acquiring skills that are imparted through conversation and apprenticeship with an expert [collins2018cognitive, ramberg1998fostering].
@@
Among several conversational goals are those of establishing a relationship between counterparts. This is an incidental goal that in human-human interaction is addressed in parallel with informational or narrational activities [dillard1989primaryandsecondoray, caughlin2010multiplegoals], which has been studied in human-human communication [wiemann1977explication].
@@ 5 User Aspects
Some users have preferences, special needs, or other characteristics that will make a conversational system more useful than a non-conversational one; for example, people with visual impairments, senior citizens, children, or those seeking coaching. Following principles of Universal Design [goldsmith2007universal, ostroff2011universal] or “Design for All”, many or even most such specific preferences will lead to design solutions that improve interaction for every user in the general case. Contextual requirements may benefit from a conversational approach: hands-free operation of a system or situations where a user’s visual attention is otherwise occupied are examples where a conversational system would well serve every user.
@@
Table 1: A model of health information seekers, based on pang2015conceptualising
@@
Users also differ in their cognitive needs around information access, e.g., in their tendency to seek information. For example, the health domain pang2015conceptualising categorizes information seekers based on their level of reading engagement and research tactics: a knowledge digger would be high on both, whereas a quick fact checker would be low on both. In decision-making, a similar distinction is made between users who want to make the best, most optimal decision possible (maximizers) versus those who just want a satisfactory outcome (satisfiers) that satisfies most of their needs. Maximizers usually engage in extensive product comparison and aspire to find the best possible option [Cheek_Schwartz_2016]. In earlier work on a critiquing recommender system, Frederix2021 found that maximizers experienced more guidance from a (conversational) critiquing system compared to a simple filtering tool, whereas novices did not. Both these examples show that conversational systems might be used, experienced, and thus evaluated quite differently depending on individual differences of the user. Using individual difference measures (such as a maximization scale) allows one to distinguish between users in the evaluation.
@@
Cultural differences in how a conversation proceeds, if not accommodated by a conversational system, may annoy and disturb users [kim1994cross]. For instance, users from cultures that value directness and efficiency might prefer concise, straightforward interactions, whereas those from cultures that emphasize politeness and formality might expect more elaborate and courteous exchanges. This can influence how questions are asked, responses are given, and even the pacing of the conversation. The notion of politeness in itself is a many-faceted behavioral characteristic, and its various aspects are prioritized differently across cultures.
@@ 6 Task
  • an information need and knowledge gap [dervin1992sense]
@@
  • interactive and mixed initiative [horvitz1999principles]
@@
In contrast, some tasks strongly benefit from the opportunities afforded by conversation, including an iterative process and agents interacting via turn-taking. Particularly, complex tasks [BYSTROM1995191, campbell1988task] that require clarification and refinement to support a user’s goal are well-served by an interaction between the system and the user.
@@
Figure 7: The tasks that are ideally solved by CONIAC systems. The model proposes two core dimensions of task complexity and task definiteness. (Iterative vs. exploratory search behavior representation adapted from white2009exploratory)
In this model, complex tasks can vary in definiteness, with implications for the nature of an interaction between a system and a user. A complex task that is more well-defined requires primarily an iterative interaction process aimed at refining the expression of the information need to be more focused or precise [belkin2001iterative]. A more ill-defined information need involves exploration, a process that involves learning and investigation, and examining a broader range of information—often uncovering unanticipated information. This process typically involves refining the information need itself. Here, we draw on the notion of iterative vs. exploratory search defined by white2009exploratory. This model explicitly reflects the interaction behaviors that are likely to be involved with these two types of search.
A user’s information journey in complex tasks is well-documented in the literature [marchionini2006, kuhlthau1991], highlighting how user states can evolve throughout the process. For instance, the Information Search Process (ISP) model [kuhlthau1991] demonstrates that a user’s initial perception of a task’s complexity might start as simple but can quickly become more complex due to the broad scope of relevant information. This suggests that CONIAC systems should be capable of detecting and modeling these evolving user states, optimizing their functions and responses accordingly to enhance the user experience.
Task switching is another common characteristic of human information-seeking behavior, observed in both web search engines and conversational assistants [SPINK2006264, trippas2024reevaluating]. In complex tasks, users often pursue multiple sub-goals, and the outcome of each sub-goal can influence subsequent actions or significantly impact the final goal. Therefore, it is crucial for users to track multiple tasks effectively. However, traditional information access systems typically leave the burden of managing these tasks to the users.
@@
In the following, we illustrate typical conversation information access tasks and how they relate to dimensions of complexity and definiteness. We also illustrate how users might dynamically traverse the space during the conversational interaction with the system [taylor1962, kuhlthau1991].
@@
In the context of scientific discovery, hypothesis generation involves investigating scientific literature. This often includes making connections between seemingly unrelated studies [swanson1987two]. A classic example is the “ABC” literature-based discovery model introduced by Swanson [Smalheiser2017], in which implicit shared items (“B” items) are analyzed to form links between two entities (“A”, “C”). Under this model, a user could be supported by a conversational system to move logically between different parts of the information space.
Evaluation of a hypothesis generation task is challenging as there is no truly “correct” answer that the user seeks. Notions of novelty, surprise, feasibility, promisingness, or interestingness are relevant [smalheiser2012beyond].
@@
Suppose a user with limited domain knowledge needs to replace an old washing machine: “I need to replace my washing machine and I want one that serves our family of five and is somewhat efficient”. We initially would classify this as a well-defined need of moderate complexity. Current online tools (e-commerce shops) would allow you to search products with some filtering and construct a set of alternatives, but it will be hard to understand the domain: e.g., what is a sufficient capacity in kg? What amount of kWh makes it efficient? Also, the number of features presented by the tool might bring the realization that there are other important things to consider (rinsing quality, time per cycle) and might overwhelm the user with the number of tradeoffs to make. Hence, a simple search might make the user aware that the need is not as well-defined and the task perhaps more complex than envisioned. However, a regular online shop will not provide the tools to help users understand the domain better. A conversational system, using techniques from critiquing recommenders [Frederix2021] would help the user to understand the domain and tradeoffs better: “based on your query, we have several efficient washing machines that have 8kg capacity or more that suit your family needs. Would you rather like a washing machine that rinses the clothes better and but will be less silent”. Based on this response, the user discovers some other (latent) needs and tradeoffs they were unaware of before, and the dialog will allow them to define them better iteratively. This will allow the user to learn about their actual preferences and the best matching products, and the system can provide a more adequate recommendation.
@@
Looking for health information is a natural task for CONIAC. Many individuals already use the internet for health advice, such as using symptom checkers to determine if they might have an underlying condition or whether they should consult a healthcare professional [cross2021search].
Research has examined health information-seeking behaviors, for example, considering how users may address an open-ended task such as “Imagine you are going to a party and will discuss health information with your friends. Gather enough information for an interesting discussion.” (adapted from [pang2016designing]). This is a vague but complex task, requiring exploration of a large and dynamic information space, accommodation of serendipity to access novel health concepts, and interactions between the user and the system to support the identification of information that is interesting and engaging to the user.
A recent study on a CONIAC system designed to help patients find cancer-related clinical trials indicates that these systems could make health information more accessible for individuals with limited health or computer literacy skills [bickmore2016improving].
While CONIAC has significant potential, several concerns must be addressed when implementing these systems in the health domain. One issue is that these systems may lack the necessary expertise to accurately answer all questions, which could lead to misunderstandings or misinterpretations and, consequently, incorrect responses [su2019improving]. Although this is a common challenge for all search systems, it could be particularly catastrophic with health information. Additionally, these systems often handle sensitive patient data, which requires robust safeguarding measures. Voice-only CONIAC systems might also face difficulties with speech recognition, especially when users are distressed or in noisy environments [spina2021SIGIRForum].
@@
Enterprise information access and personal information access have distinct requirements compared to traditional web search engines. These include challenges such as searching across enterprise (individual) intranets or navigating multiple internal or private data sources [Hawking2004enterprise]. Although there has been growing interest in workplace-oriented digital assistants, such as Alexa for Business or the Cortana Skills Kit for Enterprise, their adoption has been limited. However, with the increased adoption of LLM-based systems, new solutions like Microsoft Copilot and Google Gemini are emerging to enhance productivity and streamline information retrieval in enterprise settings [trippas2024what]. Microsoft Copilot is designed to assist employees by integrating with Office applications and providing contextual assistance and insights with LLMs. Similarly, Google Gemini aims to leverage artificial intelligence to improve search and information access across Google’s ecosystem. Despite these advancements, the use of these tools remains relatively low, suggesting that further development is needed to realize and understand the possible task and the system’s potential in the enterprise or personal information environment fully [trippas2024what, trippas2024reevaluating].
@@
Children regularly engage in information discovery tasks that involve exploring online resources to complete classroom assignments. Their default approach is to use web search engines, but they often face challenges: (i) difficulty in formulating their information needs into succinct keyword queries, (ii) a preference for using natural language or question-based queries, and (iii) trouble navigating Search Engine Result Pages to find relevant resources [azpiazu2017online, landoni2021right, DBLP:conf/sigir/PeraHML25]. A conversational information access system could be highly beneficial in this context. Such a system could ask clarifying questions to help children refine their information needs, provide specific answers when required, and recommend a limited set of resources for further exploration. However, a one-dimensional evaluation (e.g., the system’s ’goodness’ based on its ability to deliver the necessary information or user perception—especially, since children’s perceptions of their interactions with information systems do not always align with their actions) is insufficient [DBLP:journals/ijmms/BurkeABNKNOPTZ25, DBLP:journals/corr/abs-2405-02050]. As children are the primary stakeholders in this scenario, it is essential to consider factors like the suitability of the information provided and whether the vocabulary matches their literacy levels. This highlights the importance of explicitly considering the task at hand, the stakeholder performing the task, and the various criteria for evaluation before determining the specific metrics to be used as we reflect on the CAFE components.
@@
Let us suppose that a user starts interacting with a CONIAC system as follows: “The next two and a half hours, I will be working on something that takes up most of my attention, gaze, and hands. I want to have something entertaining in my headphones during that time. I am interested in some general topics: personal aviation, allergy medication, history of Central Asia, and I like yacht rock and string quartets” (the last part is presumably known to the system). The system can suggest some listening material to the user, excerpts from audiobooks and lectures, recent newscasts with local news, and playlists the user has saved in the library. The user can then respond by ranking the suggestions, helping the system to sort the items that will be played in the session. The system can verify if those preferences are generally true or specific to this session.
@@
Let us consider the following scenario. The users asks a system about interesting sights to see in Saarland and bike paths. The user gets some tips, but they do not seem to be geared toward her tastes, so the user elaborates that she wants bike paths for mountain biking and that she will be camping. The system responds but delivers the same content. The user then asks for tips on Limburg, an area she has already vacationed in and know well. The user realizes that the knowledge of this system is superficial while it was initially presented as a useful resource, and she will now turn to another system. (If not conversational, the user would have had to work hard to redo the query for another area).
@@
Figure 8: Visualization of possible tasks for information access, demonstrating different information access and activity levels [white2024tasks, trippas2024adapting]. The diagram highlights the progression from simple information finding to more complex tasks like learning, investigating, critiquing, evaluating, extracting, and creating. The new frontier with Generative Information Retrieval (marked by dashes) indicates that these systems can enable advanced tasks such as critiquing and evaluating, expanding beyond the traditional search frontier also applicable to CONIAC tasks.